40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Nov 7 14:39:49 2013, Jenne, Update, LSC, Lock Acquisition Game Plan LockAcquisition_7Nov2013.pdf
    Reply  Fri Nov 8 18:12:21 2013, Jenne, Update, LSC, PRFPMI: Not crossing any resonances 2f_resonances_Xarm_CARMsweep.pngAllModFreq_resonances_Xarm_CARMsweep.png
    Reply  Tue Nov 12 16:49:22 2013, Jenne, Update, LSC, Xend QPD and Whitening board pulled IMG_1262.JPGIMG_1260.JPGIMG_1264.JPGIMG_1276.JPG
       Reply  Wed Nov 20 17:05:15 2013, Jenne, Update, LSC, Xend QPD and Whitening board replaced 
       Reply  Wed Dec 18 00:26:15 2013, Jenne, Update, LSC, Xend QPD schematic investigation 
          Reply  Wed Dec 18 11:32:34 2013, rana, Update, LSC, Xend QPD schematic investigation 
    Reply  Wed Nov 13 01:35:40 2013, Jenne, Update, LSC, PRM motion causing trouble? xarmRIN_vs_CARM.pngASportOverlap_moveGaussianBeams.png
       Reply  Wed Nov 13 09:31:15 2013, Gabriele, Update, LSC, PRM motion causing trouble? 
          Reply  Wed Nov 13 19:22:58 2013, Jenne, Update, LSC, PRM motion correlated to intracavity power POPDC_sensors_30oct2013.pngAOMmodulated_POPDC_OSEM.pdf
             Reply  Thu Nov 14 02:50:43 2013, Jenne, Update, LSC, PRM oplev measured and modeled TF PRM_OL_TF_model.pngPRM_OLPIT_TF.pdfPRM_OLYAW_TF.pdf
                Reply  Mon Nov 18 21:02:54 2013, Jenne, Update, LSC, PRM oplev measured and modeled TF PRM_OLG_NewOld.pngPRM_oplev_improvement.pdf
                   Reply  Mon Nov 18 21:40:24 2013, Koji, Update, LSC, PRM oplev measured and modeled TF 
                      Reply  Tue Nov 19 14:47:44 2013, Jenne, Update, LSC, PRM oplev measured and modeled TF 
             Reply  Thu Nov 14 02:55:26 2013, rana, Update, SUS, PRM motion correlated to intracavity power 
                Reply  Thu Nov 14 11:41:19 2013, Steve, Update, SUS, PRM sensors effected by IR IRon_off.png
                   Reply  Thu Nov 14 14:35:12 2013, Steve, Update, SUS, IR effect on MC and PRM sensors MC_PRM_IReffect.png
                      Reply  Fri Dec 6 10:03:07 2013, Steve, Update, SUS, IR effect on MC sensors only 10minPMCnotLocked.png6Dec2013.png
Message ID: 9360     Entry time: Thu Nov 7 14:39:49 2013     Reply to this: 9362   9367   9371
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Lock Acquisition Game Plan 

Between the 40m meeting, and chatting with Gabriele, there was lots of talking yesterday about our 40m Lock Acquisition game plan. 

From those talks, here is my current understanding of the plan, in a Ward-style cartoon:

 LockAcquisition_7Nov2013.pdf(This is a 2 page document - description of steps is on 2nd page)

If you look closely, you will notice that there are several places that I have used "?" rather than numbers, to indicate what RFPD signal we should be using.  To fill these in, I need to look at some more simulations, and think more carefully about what signals exist at what ports, and what SNR we have at each of those ports.

Also, while the overall scale of the arm power plot is correct, the power level at each step is totally arbitrary right now, and should just be taken to mean places (in time) where the CARM offset is reduced a little more.


There are several things at this point that we know we need to look into:

* POP 22/110 PD and filtering electronics should be switched to a broadband PD, rather than the Thorlabs PD + Miniciruits filters. (Hardware)

* Whitening for the transmission QPDs needs to be thought about more carefully. (Calculation, then hardware)

* Chose a good SNR REFL DC signal, which may or may not be from the PD we are currently using (I think it's the DC of REFL11, but I'll have to check). (Calculation)

* For DRMI locking, what is the size of the SRCL error signal at AS55, AS165, and the REFL ports?  Do we need to lock with AS port, and then switch over to a REFL 3f port, to make acquisition easier?  (Simulation)

* Similarly, I want to make the equivalent of Figure 3 of T1000294, with our 40m parameters. (Simulation)

* To set the phase of AS110, simulate the demod phase of AS110 in both DRMI and SRMI cases.  If no (significant) change, maybe we can set the phase in the real system by misaligning the PRM, and watching the SRMI flash.  (Simulation)

* Simulate an arm sweep, up to many orders of the sidebands, to see how close to the carrier resonance any sideband resonances might be.  If something like the 4th order sideband resonates, and then beats with a 1st order sideband, is that signal big enough to disturb our 3f locking of the PRMI / DRMI?  We want to be holding the arms off resonance with ALS closer to the carrier than any "important" sideband resonances (where the definition of "important" is still undetermined).  (Simulation)

* Check if we can hand DARM from the DC transmission signals to the final RF signal while we still have a large CARM offset. Is there a point where the CARM offset is too large, and we must be still using the DC signals? (Simulation)

* At what arm power level can we transition from ALS to IR DC transmission signals for the individual arms? (Simulation)

* Still need to finish calculating what could be causing our big arm power fluctuations (Test mass angular motion?  PRM angular motion? ALS noise?) (Calculation)


Replys, and comments are welcome, particularly to help me understand where I may have (likely did) go wrong in drawing my cartoon. 

ELOG V3.1.3-